:: State ?(1), ?(2), and ?(3)
:: State the Inductive Hypothesis you would use if you were proving the statement was true for all positive

integers using mathematical induction
:: State the “Left Hand Side” of the statement in the ? = ? + 1 case

4. Σ- (-3) = 2**) 4. 2n+1+(-1)* 3:27 i=1

3. 12 – 22 + 32 – ... + (-1)n-1n2 = (–1)n-1 n(n+1)
Answer #1

3: 860=1 ?- 94+39-__ _ + (-)ima = 6in (hu) L48 = P(1)= 12 -1 I PW = $1972!1+2 = 1 = R HR PC)=12-22 = -3 || 9(2) = (-19224.4.LH: 8(x+1)= Eink! K (K) & Cik Citr vom 0} = fjk (k+) Fş - (ku)] = (-) + (xt) (-4-2) 11.15 = P(x+) = (-10% (kx) (x+2) RAD = Pnul - ). ) + (-) . PEL) = +59 +49 = ++-3 | Phn) - Saint C 3.2m N- Note: given question is wrong. the value of from jo to veriIL HS = RHS So, 1C1) is due for nal. Step2in het for n=k, f(n) is due. so, P(K)= dk (-1)} = 2 + + (-1)* 3.2* = K+1 Steps: Now

Note:just play with maths & best of luck for your bright future .

WE ARE ALWAYS PRESENT HERE IN YOUR SERVICE.

THANK YOU!